6-27. Computing Straight-Line and Double-Declining-Balance Depreciation On January 2, Reed Company purchases a laser cutting machine for use in fabrication of a part for one of its key products. The machine cost $75,000, and its estimated useful life is five years, after which the expected salvage value is $5,000. For both parts a and b below: (1) Compute depreciation expense for each year of the machines five-year useful life under that depreciation method. (2) Use the financial statements effects template to show the effect of depreciation for the first year only for that method. (When equipment is used exclusively in the manufacturing process, the depreciation is more accurately recorded as part of cost of goods sold and not as depreciation expense.) a. Straight-line b. Double-declining-balance
Answer #1
1) a) Straight line
Depreciation for each year=(Purchase value-salvage value)/life
=(75000-5000)/5=$14,000
b) double declining balance
Depreciation rate=(100/5)*2=40%
Year Opening Book value Depreciation @ 40% Closing book value
1 $        75,000 $         30,000 $        45,000
2 $        45,000 $         18,000 $        27,000
3 $        27,000 $         10,800 $        16,200
4 $        16,200 $            6,480 $          9,720
5 $          9,720 $            4,720 $          5,000
2)
Balance sheet (Double declining)
Transaction Current Asset + Asset= Liabilities+ Stockholder's equity
$       (30,000) $       (30,000)
Balance sheet (Straight line)
Transaction Current Asset + Asset= Liabilities+ Stockholder's equity
$       (14,000) $       (14,000)